OpenStudy (anonymous):

Which property is demonstrated by this equation? 3(1 + 2) = (1 + 2)3

OpenStudy (zzr0ck3r):

it might just be easier to look up all the properties. You are asking about definitions....they are easy to look up/

OpenStudy (anonymous):

k then.

OpenStudy (zzr0ck3r):

do you still not know?

OpenStudy (zzr0ck3r):

\(ab=ba\) commutative property of multiplication \(a+b=b+a\) commutative property of addition \((ab)c=a(bc)\) associative property of multiplication \((a+b)+c=a+(b+c)\) associative property of addition

OpenStudy (zzr0ck3r):

note that this looks like the third one, but its the first one